Show steps and use formulas. Do not use the table please!!!

The Seminole Production Company is analyzing the investment in a new line of business machines. The initial outlay required is $28 million. The net cash flows expected from the investment are as follows:

Year Net Cash flow (Million)
1 $8
2 12
3 19
4 24
5 21
6 20
7 15

The firm’s cost of capital (used for projects of average risk) is 14 percent. Use Table II to answer the questions below. Enter your answers in millions. For example, an answer of $1.211 million should be entered as 1.211, not 1,211,000. Round your answers to three decimal places.

  1. Compute the net present value of this project assuming it possesses average risk.

    $    million

    At a 14% rate, the project is -Select-acceptableunacceptableItem 2 .

  2. Because of the risk inherent in this type of investment, Seminole has decided to employ the certainty equivalent approach. After considerable discussion, management has agreed to apply the following certainty equivalents to the project’s cash flows:

    Year αt
    0 1.00
    1 0.95
    2 0.90
    3 0.80
    4 0.75
    5 0.65
    6 0.50
    7 0.40

    If the risk-free rate is 9 percent, compute the project’s certainty equivalent net present value.

    $    million

  3. On the basis of the certainty equivalent analysis, should the project be accepted?

    The project is -Select-acceptableunacceptableItem 4 .
Answer #1

a)

NPV is calculated using NPV function in excel

А B с D 1 Year Net cash flow in millions N 3 0 -28 1 8 4 5 12 N LD 3 19 7 4 24 8 Ол 21 9 6 20 10 7 15 11 12 NPV 41.299 C3+NPV

Calculating NPV using formula = -28 + (8/(1.14^1)) + (12/(1.14^2)) + (19/(1.14^3)) + (24/(1.14^4)) + (21/(1.14^5)) + (20/(1.14^6)) + (15/(1.14^7))

Net present value of this project = $41.299 million

Since the NPV is positive, the project is acceptable

b)

А B E F 1 2. 3 Certainty equivalent cash-flows -28 7.6 4 5 10.8 6 Year Net cash flow in millions certainty 0 -28 1.00 1 8 0.9

A B D E Year Certainty Net cash flow in millions -28 0 1 1 2 3 4 5 6 8 1 2 12 3 19 Certainty equivalent cash-flows =D3*C3 =D4

Calculating NPV using formula = -28 + (7.6/(1.09^1)) + (10.8/(1.09^2)) + (15.2/(1.09^3)) + (18/(1.09^4)) + (13.65/(1.09^5)) + (10/(1.09^6)) + (6/(1.09^7))

Project’s certainty equivalent net present value = $30.668 million

Since the NPV is positive, the project is acceptable